# Three Heads Are Better Than One: Complementary Experts for Long-Tailed Semi-supervised Learning

### v1

@@ Abstract
We address the challenging problem of Long-Tailed Semi-Supervised Learning (LTSSL) where labeled data exhibit imbalanced class distribution and unlabeled data follow an unknown distribution. Unlike in balanced SSL, the generated pseudo-labels are skewed towards head classes, intensifying the training bias. Such a phenomenon is even amplified as more unlabeled data will be mislabeled as head classes when the class distribution of labeled and unlabeled datasets are mismatched. To solve this problem, we propose a novel method named ComPlementary Experts (CPE). Specifically, we train multiple experts to model various class distributions, each of them yielding high-quality pseudo-labels within one form of class distribution. Besides, we introduce Classwise Batch Normalization for CPE to avoid performance degradation caused by feature distribution mismatch between head and non-head classes. CPE achieves state-of-the-art performances on CIFAR-10-LT, CIFAR-100-LT, and STL-10-LT dataset benchmarks. For instance, on CIFAR-10-LT, CPE improves test accuracy by over $>$2.22% compared to baselines. Code is available at https://github.com/machengcheng2016/CPE-LTSSL.
@@ 2.2 Long-tailed learning
Since the class distribution of the training set is usually imbalanced, existing algorithms in long-tailed learning attempt to improve the generalization ability of DNN towards the medium and tail classes. Generally, the framework of existing methods can be summarized into four categories: Re-sampling (Chawla et al. 2002; Buda, Maki, and Mazurowski 2018) conducts under-sampling on head classes or over-sampling on tail classes to balance the training data; Re-weighting (Khan et al. 2017; Cui et al. 2019) adjusts the class weight or sample weight to maintain a balanced training process; Ensemble learning (Wang et al. 2020; Cai, Wang, and Hwang 2021; Zhang et al. 2022) is based on multiple experts (classifier heads) to enhance the representation learning; Loss modification (Cao et al. 2019; Hong et al. 2021) modifies the logit value by margins in either the training or inference stage and the typical work “logit adjustment” (Menon et al. 2020) which has been proved as the Bayes-optimal solution of the long-tailed problem. Nonetheless, the current approaches are predominantly tailored for the fully-supervised context, rendering them unsuitable for semi-supervised learning where the class distribution is unspecified.
@@ A.1 Logit adjustment intensities for multi-experts
The hyper-parameters $(\tau_{1},\tau_{2},\tau_{3})$ decide the logit adjustment intensities for multi-experts to model the three different class distributions. For example, a small set of $(\tau_{1},\tau_{2},\tau_{3})$ can only cover the ((((consistent, less consistent, uniform)normal-))) cases, but fail to cover the inverse case. We test on a set of $(\tau_{1},\tau_{2},\tau_{3})$ with the number of experts being 2 or 3 in the inverse case, and we show the corresponding performances in Table 5 and Fig. 6. We can see that the test accuracy on tail classes significantly drops if the third expert is removed, or when $\tau_{2}$ and $\tau_{3}$ are not large enough. Such results emphasize the significance of multi-expert model where each expert can cover one specific class distribution and generate high-quality pseudo-labels in each corresponding case.
@@
A.2 Use all three experts or just $E_{2}$?
The three heads are experts in skewed/balanced/inversely-skewed distributions. Since the test set distribution is usually balanced, CPE only utilizes $E_{2}$ during testing. In fact, we tried to utilize the average of three predictions for testing (see Table 6), and the performance slightly drops. The reason is that $E_{1}$ and $E_{3}$ yield large probabilities on head and tail classes, respectively, which hurt the performance on medium classes.

### v2

@@ Abstract
We address the challenging problem of Long-Tailed Semi-Supervised Learning (LTSSL) where labeled data exhibit imbalanced class distribution and unlabeled data follow an unknown distribution. Unlike in balanced SSL, the generated pseudo-labels are skewed towards head classes, intensifying the training bias. Such a phenomenon is even amplified as more unlabeled data will be mislabeled as head classes when the class distribution of labeled and unlabeled datasets are mismatched. To solve this problem, we propose a novel method named ComPlementary Experts (CPE). Specifically, we train multiple experts to model various class distributions, each of them yielding high-quality pseudo-labels within one form of class distribution. Besides, we introduce Classwise Batch Normalization for CPE to avoid performance degradation caused by feature distribution mismatch between head and non-head classes. CPE achieves state-of-the-art performances on CIFAR-10-LT, CIFAR-100-LT, and STL-10-LT dataset benchmarks. For instance, on CIFAR-10-LT, CPE improves test accuracy by over 2.22% compared to baselines. Code is available at https://github.com/machengcheng2016/CPE-LTSSL.
@@ 2.2 Long-tailed learning
Since the class distribution of the training set is usually imbalanced, existing algorithms in long-tailed learning attempt to improve the generalization ability of DNN towards the medium and tail classes. Generally, the framework of existing methods can be summarized into four categories: Re-sampling (Chawla et al. 2002; Buda, Maki, and Mazurowski 2018) conducts under-sampling on head classes or over-sampling on tail classes to balance the training data; Re-weighting (Khan et al. 2017; Cui et al. 2019) adjusts the class weight or sample weight to maintain a balanced training process; Ensemble learning (Wang et al. 2020; Cai, Wang, and Hwang 2021; Zhang et al. 2022; Aimar et al. 2023) is based on multiple experts (classifier heads) to enhance the representation learning; Loss modification (Cao et al. 2019; Hong et al. 2021) modifies the logit value by margins in either the training or inference stage and the typical work “logit adjustment” (Menon et al. 2020) which has been proved as the Bayes-optimal solution of the long-tailed problem. Nonetheless, the current approaches are predominantly tailored for the fully-supervised context, rendering them unsuitable for semi-supervised learning where the class distribution is unspecified.
@@ A.1 Logit adjustment intensities for multi-experts
The hyper-parameters $(\tau_{1},\tau_{2},\tau_{3})$ decide the logit adjustment intensities for multi-experts to model the three different class distributions. For example, a small set of $(\tau_{1},\tau_{2},\tau_{3})$ can only cover the ((((consistent, less consistent, uniform)))) cases, but fail to cover the inverse case. We test on a set of $(\tau_{1},\tau_{2},\tau_{3})$ with the number of experts being 2 or 3 in the inverse case, and we show the corresponding performances in Table 5 and Fig. 6. We can see that the test accuracy on tail classes significantly drops if the third expert is removed, or when $\tau_{2}$ and $\tau_{3}$ are not large enough. Such results emphasize the significance of multi-expert model where each expert can cover one specific class distribution and generate high-quality pseudo-labels in each corresponding case.
@@
A.2 Use all three experts or just $E_{2}$ for inference?
The three heads are experts in skewed/balanced/inversely-skewed distributions. Since the test set distribution is usually balanced, CPE only utilizes $E_{2}$ during testing. In fact, we tried to utilize the average of three predictions for testing (see Table 6), and the performance slightly drops. The reason is that $E_{1}$ and $E_{3}$ yield too large probabilities on head and tail classes, respectively, and hurt the performance on medium classes.
@@ C.3 Does CPE work with other distributions?
Appendix D Connection to recent methods
Also based on multiple experts, SADE (Zhang et al. 2022) learns towards three simulated class distributions for test-agnostic long-tailed setting. At test time, SADE aggregates the softmax predictions of all experts in a self-supervised manner for a better accuracy. Afterwards, BalPoE (Aimar et al. 2023) improves SADE and is proved to achieve calibrated and unbiased predictions. With respect to each training sample, we note that the predicted softmax scores of all experts are always learned towards a same class label. In contrast, our CPE method allows each expert to learn towards each guessing on unlabeled sample, thus one of the three experts can always yield high-quality pseudo-labels.